OpenStudy (anonymous):

Can anyone help with these two last questions of my homework because I don't understand!!? Solve for x..... 2. (1/2) + (4/2x) = (x+4/10) 9. (5/x^2-4) + (2/x) = (2/x-2)

OpenStudy (asnaseer):

do you know how to get rid of the fractions here?

OpenStudy (anonymous):

no i don't. I don't know how to add fractions at all...stupid isn't!!

OpenStudy (asnaseer):

don't put yourself down - I'm sure you are capable of learning. we've all been there :-)

OpenStudy (asnaseer):

ok, first let me confirm the equation....

OpenStudy (asnaseer):

is the first one:\[\frac{1}{2}+\frac{4}{2x}=\frac{x+4}{10}\]

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok - now what you need to do is see what to multiply both sides by in order to get rid of the fractions. we can do this one step at a time so that you can understand the process.

OpenStudy (anonymous):

by 2 !?!

OpenStudy (asnaseer):

so 1st I see a 2 in the denominator of the first fraction on the left, so lets multiply both sides by 2 to get:

OpenStudy (asnaseer):

yes - you DO understand! :-)

OpenStudy (asnaseer):

multiplying by 2 we get:\[2*\frac{1}{2}+2*\frac{4}{2x}=2*\frac{x+4}{10}\]

OpenStudy (anonymous):

but don't you want the denominators the same. wouldn't you multiply by 2x

OpenStudy (asnaseer):

now we cancel out common factors to get:\[\cancel{2}*\frac{1}{\cancel{2}}+\cancel{2}*\frac{4}{\cancel{2}x}=\cancel{2}*\frac{x+4}{\cancel{10}5}\]

OpenStudy (asnaseer):

I wanted to show you one step at a time, but if you are more confident, then we can go faster?

OpenStudy (anonymous):

no its okay i was just curious

OpenStudy (asnaseer):

ok, so we end up with:\[1+\frac{4}{x}=\frac{x+4}{5}\]

OpenStudy (asnaseer):

next, as you spotted already, we can multiply both sides by x to get:\[x+4=\frac{x(x+4)}{5}\]

OpenStudy (asnaseer):

finally, we multiply both sides by 5 to get rid of the final fraction to get:\[5x+20=x(x+4)\]

OpenStudy (asnaseer):

so now we are left with a quadratic equation. do you know how to solve from here?

OpenStudy (anonymous):

yes two questions though.... 1. why did you multiply both side by x? 2. in the quadratic formula would it x^2 or 2x?

OpenStudy (asnaseer):

ok, when we got to this step above:\[1+\frac{4}{x}=\frac{x+4}{5}\]I noticed that we still have a fraction on the left hand side, and that this fraction has an "x" in the denominator. so we multiply both sides by "x" to get rid of this denominator.

OpenStudy (anonymous):

and that canceled out the 1?

OpenStudy (asnaseer):

not cancelled out - 1 times "x" is equal to "x"

OpenStudy (anonymous):

ohh okay

OpenStudy (asnaseer):

secondly, quadratic equations involve \(x^2\). so the final equation we got was:\[5x+20=x(x+4)\]which expands to:\[5x+20=x^2+4x\]and finally to:\[x^2-x-20=0\]

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (asnaseer):

hopefully you can do the 2nd equation yourself. do you want to try it out yourself now and show me your steps?

OpenStudy (anonymous):

ill try it out but im late for work so i have to go thank you a lot though . it makes a lot more since to me know.

OpenStudy (asnaseer):

np - I'm glad I was able to explain it clearly enough for you to understand.